OpenStudy (anonymous):

The center of a circle is at the origin. An endpoint of a diameter of the circle is at (-3, -4). How long is the diameter of the circle?

OpenStudy (jhannybean):

This is the same kind of question as your last one.

OpenStudy (jhannybean):

Equation of circle : \(\large (x-h)^2 +(y-k)^2 =r^2 \) Center/(vertex) = \(\large (h,k)= (0,0)\) plug it into formula \(\large {(x-0)^2 +(y-0)^2 = r^2 \\ x^2 +y^2 =r^2}\) A diameter is made up of 2 radiuses. so we'd be solving for r^2. point given: \(\large (-3,-4)\) \(\large (-3)^2 +(-4)^2=r^2(\text{or} \ D)\) \(\large D =?\)

OpenStudy (anonymous):

10

OpenStudy (anonymous):

wait the answer is 5 right? because its 25 so the square root of that is 5

OpenStudy (jhannybean):

No, it would have been 25 because r^2 = D.